OpenStudy (nikki_skellington):

Sara left a bin outside in her garden to collect rainwater. She notices that 1 over 8 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

OpenStudy (mathstudent55):

The full bin contains x amount of water. Using x, how much water is in the bin now?

OpenStudy (nikki_skellington):

1/8 × 2/3 gal = x be a good equation and the answer is 1/12 right

OpenStudy (mathstudent55):

Don't go so fast. Just answer a question at a time. Incidentally, that equation is incorrect.

OpenStudy (mathstudent55):

Once again, the full bin contains x gallons. It is 2/3 full. Using x, can you write an expression that shows how full the bin is?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The full bin can hold x gallons of water. The bin is 2/3 full. The amount of water in the bin is \(\dfrac{2}{3} x\) We are told that the amount of water in the bin is 1/8 of a gallon, so the equation is: \(\dfrac{2}{3} x = \dfrac{1}{8} \)

OpenStudy (mathstudent55):

Now you need to solve the equation.